OpenStudy (anonymous):

A stone is dropped off a cliff and hit the ground with a speed of 120 ft/s . a) what is the equation for acceleration a(t), velocity v(t), and displacement/ position s(t)? b) what is the height of the cliff ?

OpenStudy (amistre64):

dropped off would indicate no initial velocity; so v(0) = 0 v(0) = -32(0) + C = 0 when C=0 v(t) = -32t ................................. v(t) = 120 = -32t; when t = |-15/4| this tells us it fell for 15/4 seconds before hitting the ground. ................................... s(t) = -16 (15/4)^2 + Hi = 0, when Hi = 225 the initial height of hte cliff is 225 ft the eqs are: s(t) = -16 t^2 + 225 v(t) = -32t a(t) = -32